OpenStudy (anonymous):

Find the area of a sector with a central angle of and a radius of 18.8 m

OpenStudy (anonymous):

For a sector of theta radians in a circle of radius r, the area A is: A = r^2 theta / 2. When theta = 2pi / 15 and r = 18.8m: A = 18.8^2 * pi / 15 = 74.0 m^2 to 3 sig. fig. There are 2pi radians in one revolution. Thus, from the area pi r^2 of the circle, the area of a sector subtending theta at the centre is: (theta / 2pi)(pi r^2) = r^2 theta / 2.